OpenStudy (anonymous):

Find the distance between the points (11, 4) and (10, 5). what is and easy way to find this..........will fan and medal

undeadknight26 (undeadknight26):

ok to find the slope i believe u are asking is like so: y2 - y1/x2 - x1

OpenStudy (anonymous):

The distance formula is very similar to the Pythagorean theorem: \[\sqrt{(DeltaX)^{2}+(DeltaY)^{2}}\] So you have Change in X =(11-10)=1 Change in Y = (4-5)= -1 \[\sqrt{(1)^{2}+(-1)^{2}}=\sqrt{2}\] ~= 1.41

OpenStudy (anonymous):

so \/2 is the awsner

OpenStudy (anonymous):

Sqaure root